# Leading by the Nodes: A Survey of Film Industry Network Analysis and Datasets

### v1

@@ Abstract
This paper presents a comprehensive survey of network analysis research on the film industry, aiming to evaluate its emergence as a field of study and identify potential areas for further research. Many foundational network studies made use of the abundant data from the Internet Movie Database (IMDb) to test network methodologies. This survey focuses more specifically on examining research that employs network analysis to evaluate the film industry itself, revealing the social and business relationships involved in film production, distribution, and consumption. The paper adopts a classification approach based on node type and conducts network analyses of the reviewed literature, mapping citation and co-authorship networks. The findings provide insights into the structure and interconnectedness of the field, offering potential clusters of debates and shedding light on the institutional, geographical, and demographic characteristics shaping this field. In addition, this survey contributes to understanding film industry network analysis and informs researchers interested in network methods within the film industry and related cultural sectors.
@@
Social network analysis, Film industry, Node classification, Centrality
@@ 1 Introduction
In this paper, we present a survey of published research applying network analysis methods to the study of the film industry. The film industry has long been a rich source of information for network analysis. Use of the Internet Movie Database (IMDb) is prevalent in foundational network studies that require exemplary real-world applications for network theory ([5, 2, 44, 26, 31]). [19] has identified the ready availability of film industry data via the IMDb API as a key determinant in a range of defining studies. The memorable Six Degrees of Kevin Bacon game ([25]) embedded the idea of the film industry as a collaboration network in the popular imagination. For academic studies, it was [66] that used this dataset to establish a long line of work on small-world collaboration networks based on film data as a use case ([4, 55, 31]). Given the critical importance of film industry data to the theoretical development of network science, we were interested to switch perspective to understand the ways in which network science has developed a perspective of the film industry as an inherently collaborative sector. In this article, we present the first systematic survey of this area of research, in order to guide network scientists interested in analysing the film industry on how they might usefully approach this sector given what has already been done.
@@
Our overarching goal in surveying this literature is twofold: firstly, to establish what has been undertaken in this area, and secondly, to evaluate how scholarship focussed on film industry networks has evolved as an emerging field of study. We aim to provide insights for scholars interested in utilizing network methods to explore the film industry, as well as other cultural industries, by demonstrating where existing scholarship has clustered into emergent debates and approaches, and where new research might be directed. In particular, we highlight the research lines that have advanced the least, and identify gaps in the existing literature that future network analyses of the film industry could address. With these considerations in mind, we complement our survey with a network analysis of the interconnectedness of the papers that we review in this paper. Bibliometric research and work from the science of science have shown that academic fields cannot be fully understood without consideration of the consolidation of relationships among actors in the field ([30, 20]). As such, examination of the interactions among publications and authors has much to offer a survey paper such as ours, as it enables us to offer a structural perspective on how scholarship has developed in this emerging research area.
Our analysis reveal that initial film network research primarily focused on individuals, with recent years witnessing a broader scope including production companies and geographic units. Additionally, we observe a disproportionate emphasis on the US film industry in foundational literature, highlighting the need for further exploration across global film industries. Moreover, there is potential for more studies in this field to embrace open science principles, especially considering the datasets are often derived from publicly available sources with minimal ethical concerns. Emphasizing the importance of acknowledging and building upon existing literature to foster the development of the research area, there is a pressing need for increased engagement and cross-citation within the scholarly community, particularly among researchers employing advanced computational methods in film industry analysis.
@@ 2 Scope and Analytical Strategy
Drawing on the science of science, where network methods have shed light on co-authorship and citation practices shaping academic disciplines ([30]), we constructed a paper-to-paper citation matrix and a person-to-person co-authorship matrix. Citation and co-authorship networks have been studied to gain insight into academic performance, identify innovation, measure influence or impact, accelerate discovery and to describe socio-demographic inequality in scientific publication ([71, 45, 40, 61]).
For the citation network, we constructed an adjacency matrix on a paper-by-paper basis. In this matrix, the element $[i,j]$ holds a value of $1$ when paper $i$ referenced paper $j$, and 00 otherwise. This method produces a directed binary network, illustrating the citation connections among the papers. The node set is confined to the 51 papers encompassed within our review. Furthermore, we introduced node attributes like publication type, the count of citations according to Google Scholar (as of December 2022), and the specific role a node assumes within the content of the paper. In order to better understand the demographic profiles of authors within this area of scholarship, we also created a person-by-person adjacency matrix. Within this matrix, the cell $[i,j]$ value corresponds to the frequency with which authors $i$ and $j$ have collaborated as co-authors on shared publications. This results in a weighted undirected network of co-authorship ties between authors; the set of authors is determined by the list of 106 unique authors identified on the papers included in our review. Additional node attributes were added to this network. For each author, the institutional affiliation and its geographical location were extracted from the paper. In some instances, it was necessary to differentiate (using separate node attributes) between an author’s affiliation within a research center or group and their departmental affiliation. For authors whose affiliations changed across different publications, we track each listed affiliation that is different from the initial affiliation.
@@ 3 Survey of the Literature
Survey papers typically adhere to the established structure of their respective fields. However, the literature we review spans multiple disciplines, including film industry studies, business studies, cultural policy studies, sociology, and computational sciences. These disciplines and research fields are structured very differently which complicates the task of conducting a coherent survey. One solution would be to categorize the papers by field or discipline. However, discerning the appropriate disciplinary classification for certain papers is not a straightforward process.
To address these challenges, we devised a novel solution involving the classification of papers based on node types. Through systematic examination, we found that scholarship exploring networks of individuals involved in the filmmaking process is often substantively and methodologically different from research focused on connections between films, organizations, events, or geographical units. Although a small number of research papers investigate distinct networks with different node types, to our knowledge there has been no significant study to date that considers the interdependencies among multiple types of nodes within film networks. We augment our survey of the literature with a network analysis of the papers we review.
@@ 3.1.5 Gatekeeping
Network analysis has proved especially valuable to several authors exploring the importance of gatekeeping as both a critical and as an employment practice in the film industry. [9] test the hypothesis that industry peers who act as evaluators tend to reinforce dominant social beliefs and norms, and are more likely to acknowledge core members over peripheral network members. Using Hollywood data from 1992 to 2004 they constructed film collaboration networks and analyzed the bipartite affiliation networks by creating an individual-by-movie matrix. The authors also created an “accumulative relational profile” by using moving windows of three-year intervals. Their findings suggest that social relationships play a crucial role in determining how talent is perceived and acknowledged. [10] extend this study by drawing on field theory to hypothesize that peers in the film industry favor creators at the core of the collaboration network, while critics may prefer those at the periphery. These hypotheses are tested using a dataset of over 12,000 crew and cast members who worked on movies distributed by eight major studios in the United States from 1992 to 2004. The dataset includes information on directors, writers, leading and supporting actors/actresses, editors, cinematographers, and production designers, sourced from IMDb. They use a discrete choice model to analyze award decisions by organizations and estimate the choice set of eligible professionals. Collaboration networks are created based on person-film affiliations and analyzed using a core-periphery algorithm. Node classification as core or periphery is then used as a predictor in the models. Results show that being in the periphery negatively affects peer recognition but has a slight positive association with recognition by critics, controlling for other factors.
@@ 3.4.2 Reward Networks
4 A Network Analysis of Papers on Film Industry Networks
To analyze research structures in specific subject areas, we look into the co-authorship and citation networks produced by the literature we have reviewed, including the contributions of specific institutions or places to this scholarship network. Our first goal is to understand the structure and composition of the co-authorship network (Figure 2) in order to explore the profiles and collaborative relationships of those that have contributed to the field of film industry network analysis to date. We observe there is little connectivity in the co-authorship network, which has 106 nodes and 37 components, the largest of which has just 10 nodes. This relates to an interesting finding, that authors with multiple publications are inclined to collaborate with their previous colleagues, consolidating rather than expanding their connections. The components in the co-authorship network are mostly comprised of scholars based in the same country as each other, suggesting that while this emerging form of scholarship is globally dispersed, authorship collaborations tend to be nationally bounded. Authors in the US ( blue nodes) tend to have the highest international co-author rates, while countries such as Australia ( green nodes) and Brazil ( light orange nodes) have high national co-authorship rates but low international rates. Noting these tendencies, we probe further to understand how this emerging field reveals power relations and global inequalities in knowledge production. We employed Gephi software for visualizations, NetworkX for network creation, manipulation, and analysis, and utilized Pandas and NumPy for data formatting.
Figure 2: The collaborative research network with $106$ authors, color-coded by country affiliations. Node size represents author citations, link thickness indicates collaborations, and lead authors with top citations are highlighted.
A key question we want to ask of our network datasets is whether film network research adopts the geographical skew that has been noted in science of science literature ([18]). We adjust a conventional “West versus the Rest” or “Global North versus Global South” division by testing instead how the emerging literature of film network research is positioned in terms of the more nuanced core-periphery description of Majority World Countries (most of Africa, Asia, South and Central America and the Caribbean, and the Middle East) and Minority World Countries (primarily North America, Europe, and Australasia) ([12, 3]). To answer this question, we derive a country collaboration network from the co-authorship network, based on the institutional affiliation of the authors at the time of publication (Figure 3). In this network, a node represents a country, and the thickness of the ties between these nodes represents the number of times an author affiliated with a research institution in country $i$ co-authored a film networks paper with an author affiliated with an institution in country $j$. We find that scholars in the USA have dominated this emerging research area – nearly $30\%$ of the publications we reviewed were led by US-based researchers, followed by Germany and Italy with around $14\%$ each, and China with almost $10\%$. The USA holds the most central position in the country collaboration network in terms of its degree centrality, with US-based scholars collaborating with co-authors from institutions in Europe, Asia and South America. We also found that the publications by researchers based in the USA, Italy and Germany received the most citations by far. Interestingly, we found that papers looking at networks of individual people in the film industry were largely published by scholars in North America and Europe, while scholars in Asia were more likely to investigate networks of collective entities such as companies or places. We also note that as this field emerges, an increasing number of papers from the Majority World are being published.
Figure 3: Mapping global film industry research collaboration: Nodes represent countries, linked when authors co-author. Node size indicates degree centrality, link thickness reflects collaboration intensity, and color signifies continent affiliation.
Table 4 further details the publication distribution at regional (among institutions between the same province/state) and national (among institutions between different provinces/states of a country) levels. We observe that, similar to the international collaboration, the USA, Germany, and Italy lead in regional and national collaborations as well. Undoubtedly, publications with the most citations (i.e., above 500 citations as of March 18, 2023) are from these three countries and in the same order. The table also shows the gender distribution of the lead authors. While publications led by men outnumber those led by women and gender minorities (WGM) in the USA and Italy, we see a numerical balance of lead authors in Germany. Looking at publications based on the node type, nearly $78.6\%$ of the papers led by USA institutions focus on people as nodes, followed by Italy with $71.4\%$, and Germany with $50\%$. Interestingly, if we look at the number of publications, nearly $63.6\%$, $50\%$, and $80\%$ of the output of articles from these three countries are led by men authors, respectively. In the film node category, these three countries contribute equally. However, $80\%$ of the publications led by Chinese institutions focus on companies as the constituent nodes in their analysis, all led by WGM researchers. Moreover, WGM researchers tend to lead the research in China. While the main focus on Minority World publications examines individuals as nodes, newer research from Majority Countries demonstrates greater diversity of subject area as well as authorship gender and location. Table 4 also summarizes the number of institutions involved in research outcomes for each jurisdiction. There are 61 institutions involved in total, but since Ireland and Switzerland do not lead in any of the publications, they are excluded from the table. However, out of all the 58 institutions that have lead authors, $34\%$ are located in Europe, $28\%$ are in the USA, and $22.4\%$ in Asia. A positive correlation emerges between the citations received by research outputs from these nations and the number of institutions engaged within each respective country. This underscores the early initiation of research on this subject within the Minority World, propelled by the swift evolution of interest in the European/English language film industries. Concurrently, as the film industry in the Majority World undergoes its own evolution, film networks have attracted increasing scrutiny in recent times.
Table 4: Global network statistics on collaborative publications led by different countries, the gender distribution of the lead authors, and the number of institutions involved.
Publications Gender distribution
Country Regional National International Men WGM Institutions
USA 7 1 6 10 4 16
Germany 5 1 2 4 4 9
Italy 4 1 3 7 1 7
China 2 0 3 0 5 6
Australia 1 1 0 1 1 4
Brazil 2 0 0 1 1 4
South Korea 2 0 0 2 0 2
Hungary 0 0 1 1 0 2
Türkiye 2 0 0 2 0 2
Iran 1 0 0 1 0 1
Israel 1 0 0 0 1 1
Netherlands 1 0 0 1 0 1
Russia 1 0 0 1 0 1
South Africa 2 0 0 2 0 1
UK 1 0 0 1 0 1
Total: 32 4 15 34 17 58
[…]
Figure 4 shows the institutional network and the collaborations among them. The network comprises 61 institutions, 29 links, and has a lower average degree of 0.951, density of 0.016, and average clustering coefficient of 0.921. This is because each node may encompass multiple collaborations between researchers affiliated with the same institution. According to the normalized degree centrality, the most central organizations in the network are, in order of importance, New York University (USA), Universita’ di Bologna (Italy), Deakin University (Australia), RMIT University (Australia), The University of South Wales (UNSW; Australia), and the University of Technology (UTS; Australia). It is noteworthy that only two (University of Pennsylvania and UNSW) of the five prestigious institutions have engaged in collaborations with four of the universities that hold the highest degree values. Desite the sparseness of the network, we also discern a correlation between the high-degree nodes in the institutional network and the co-authorship network depicted in Figure 5. Authors with multiple publications are inclined to collaborate with their previous colleagues, contributing to the relative centrality of their respective clusters within the network.
Figure 4: The network of the 61 institutions involved in film network research, color-coded based on their country of affiliation. The node size corresponds to the degree centrality and the prestigious institutions are explicitly shown in red.
Table 5: Top six central institutions in the research of global film network.
Institution Country Degree Betweenness Closeness Eigenvector
New York University USA 0.0882 0.0012 1 0.5087
Universita’ di Bologna Italy 0.0882 0.0012 1 0.5087
Deakin University Australia 0.0588 0 1 1
RMIT Australia 0.0588 0 1 1
UNSW Australia 0.0441 0.0007 1 1
UTS Australia 0.0441 0.0007 1 1
Northwestern University USA 0.0294 0.0023 1 0.0563
[…]
00footnotetext: Centrality values were normalized in accordance with the size of the network.
Some organizations, while not prominently ranked in terms of degree centrality, can play central roles based on their betweenness or eigenvector measures. As shown in Table 5, New York University and Universita’ di Bologna appear to exhibit a higher degree of local centrality. However, their co-authors do not appear to be strongly interconnected, leading to lower eigenvector values in comparison to their Australian counterparts. The Northwestern University in the USA serves as a prominent example of an intermediary institution, effectively bridging (i.e., high betweenness) two institutions that would otherwise remain unconnected. This position bestows it with significant power and influence, as it is poised to regulate the flow of information and knowledge between these distinct universities. Within this network, closeness centrality values are uniformly distributed among the highly connected institutions, signifying that most organizations can efficiently access and share information.
Figure 5: Citation network of 51 publications: Nodes color-coded by lead author’s country. Node size represents (a) in-degree (citations received) and (b) out-degree (citations given) within the network.
The citation network to explore how the contributions to the literature have engaged with existing work and served as a reference point for future publications. By examining how these 51 publications on film network analysis cite and are cited by one another, we can reveal the influence and interconnectedness of ideas within this specific field. In the citation network given in Figure 5, in-degree centrality measures the number of citations a publication receives, highlighting its perceived importance and impact, while out-degree centrality quantifies the citations a publication makes, indicating its role in disseminating knowledge. These measures are heavily linked to time and sequence, as a paper can only realistically cite papers that have come before it, such that we would expect higher in-degree scores for older papers and higher out-degree scores for newer papers.
Figure 6: Citation network of 51 publications: Nodes color-coded by lead authors from Majority (red) and Minority (blue) Worlds. Node size represents (a) in-degree (citations received) and (b) out-degree (citations given) within the network.
The citation network is much more cohesive than the co-authorship network, suggesting that the papers we review have crystallized into a relatively well-connected set of debates over time. The network depictions shown in Figure 5a do, however, highlight some interesting patterns in who has done the work of stitching this emerging literature together. Papers led by US-based authors are notably prominent among the in-degree scores, suggesting that they have been seen as particularly important and influential by scholars using networks to study the film industry. To an extent, this can be explained by the publication dates of these papers, several of which are early, foundational studies in the area. However, an interesting dimension to the pattern is that many of the papers citing the US-based papers are from other parts of the world. In fact, when we examine out-degree scores in general (Figure 5b), we see a very different picture: scholars based in institutions outside of the US and Europe are much more likely to reference earlier work than their counterparts based in the global academic elite.
Two other interesting patterns emerge from the citation network. First, we note that many of the papers that are dotted around the edge of the citation network (indicating that they have fewer links with the other papers in either direction) are contributions from fields that do not specialize in research on the creative industries (such as computer science, physics and mathematics). By contrast, the network core is dominated by sociological papers. Moreover, the peripheral papers do not connect themselves much to prior work, despite being among the more recent publications in our review. Upon closer inspection, many of these papers are also conference papers and other non-journal article publication types. It seems that it is less common for this type of contribution to concern itself with what others have done before, instead presenting itself as a novel approach. Second, many of the papers that are embedded in the heart of the citation web are papers concerned with networks of people working in the film industry. This is related to the first point in that the interest in interpersonal networks in the film industry is largely sociological, but it indicates that there is more of an emergent discourse in this type of network study than the others. This discrepancy may guide future scholars wishing to contribute further to this area of research, as it highlights that more work is needed to build theory from the sporadic evidence provided by networks of other types of entities.
Table 6: Top six central lead authors in the research of global film network.
Lead Author Country Degree Betweenness Closeness Eigenvector
Deb Verhoeven Australia 0.0762 0.0011 1 1
Giacomo Negro Italy 0.0571 0.0035 0.75 0.1319
Simone Ferriani Italy 0.0571 0.0043 0.833 0.1386
Jonas S. Krauss Germany 0.0476 0.0001 1 0.1987
Max Wasserman USA 0.0476 0 1 0.3048
Gino Cattani USA 0.0381 0.0005 0.6428 0.116
[…]
00footnotetext: Centrality values were normalized in accordance with the size of the network.
Table 6 compares the lead researchers having the highest degree centrality measure in the network of Figure 5. According to this centrality measure, the most central people in the network are Deb Verhoeven (Australia), Giacomo Negro (Italy), Simone Ferriani (Italy), Jonas S. Krauss (Germany), Max Wasserman (USA), Gino Cattani (USA), and Xin-Jie Zhang (China). In terms of serving as an intermediary linking two or more people, S. Ferriani seems to have a higher betweenness centrality as he has collaborated with both G. Negro and G. Cattani in multiple research publications. This translates to likely being able to control the information and knowledge flow between these two research teams. In terms of how short the shortest paths are from a given author to the others, we see that even though the cluster involving G. Negro, G. Cattani, and S. Ferriani is highly modular, they are not as tight-knit as the other lead authors within their cluster. Similar to degree centrality, eigenvector centrality assesses an author’s impact by considering the quantity of its connections to other authors within the network. However, eigenvector centrality extends this concept by factoring in not only a node’s connections but also the quality of those connections, including the connections of its connections, and so forth throughout the network. The table clearly reveals that, despite having low degree and betweenness values, Max Wasserman is linked to authors who, in turn, possess high degrees of centrality. On the contrary, Gino Cattahni, who has far more publications, exhibits the lowest eigencentrality because he has co-authored with a relatively smaller group of people.
Figure 7: Country-wise research contributions based on node types. Nodes are sorted based on the total citations received, while edge thickness reflects the number of research papers in each node category.
Figure 7 complements Figure 1 by illustrating the contribution of papers in each node type led by individual countries. Notably, approximately $58\%$ of the papers concentrate on people, $16\%$ on films, $11\%$ on companies, $9\%$ on geographic units, and the remaining $6\%$ on other events. The Minority World demonstrates significant activity across all five categories, particularly emphasizing on networks related to people, films, and companies, which accounts for their higher citation rates. Another intriguing observation within Majority World countries is that China exhibits greater diversity in this research domain, contrasting with most other countries in this category, which majorly concentrate on people within film networks.
Figure 8: Citation network of 51 publications: Nodes color-coded by type people (brown square), films (green square), company (blue square), geographic units (black square), and other events (orange square). Node size represents (a) in-degree (citations received) and (b) out-degree (citations given) within the network. Pink triangles represent hybrid papers that fall in more than one node category.
Finally, Figure 8 shows the cited (in-degree) and citing (out-degree) papers based on the node type investigated. In Figure 8a, early research mainly focused on people networks in the film industry, while research on other type of network nodes is relatively recent. In Figure 8b, papers that focus on networks based on node types other than people do not tend to cite the fundamental papers (the larger square nodes), but rather more technical papers beyond the scope of film industry networks or papers analyzing networks in which people are not the constituents.
In the next section we highlight the key findings of our survey and their implications for future network-based research in this area.

### v2

@@ Abstract
This paper presents a comprehensive survey of network analysis research on the film industry, aiming to evaluate its emergence as a field of study and identify potential areas for further research. Many foundational network studies made use of the abundant data from the Internet Movie Database (IMDb) to test network methodologies. This survey focuses more specifically on examining research that employs network analysis to evaluate the film industry itself, revealing the social and business relationships involved in film production, distribution, and consumption. The paper adopts a classification approach based on node type and summarises the key contributions in relation to each. The review provides insights into the structure and interconnectedness of the field, highlighting clusters of debates and shedding light on the areas in need of further theoretical and methodological development. In addition, this survey contributes to understanding film industry network analysis and informs researchers interested in network methods within the film industry and related cultural sectors.
@@
Social network analysis, Film industry, Node classification, Collaboration networks, Literature survey
@@ 1 Introduction
In this paper, we present a survey of published research applying network analysis methods to the study of the film industry. The film industry has long been a rich source of information for network analysis. Use of the Internet Movie Database (IMDb) is prevalent in foundational network studies that require exemplary real-world applications for network theory ([4, 2, 38, 22, 26]). [16] has identified the ready availability of film industry data via the IMDb API as a key determinant in a range of defining studies. The memorable Six Degrees of Kevin Bacon game ([21]) embedded the idea of the film industry as a collaboration network in the popular imagination. For academic studies, it was [58] that used this same dataset to establish a long line of work on small-world collaboration networks based on film data as a use case ([3, 48, 26]). Given the critical importance of film industry data to the theoretical development of network science, we were interested to switch perspective to understand the ways in which network science has developed a perspective of the film industry as an inherently collaborative sector. In this article, we present the first systematic survey of this area of research, in order to guide network scientists interested in analysing the film industry on how they might usefully approach this sector given what has already been done.
@@
Our overarching goal in surveying this literature is twofold: firstly, to establish what has been undertaken in this area, and secondly, to evaluate how scholarship focussed on film industry networks has evolved as an emerging field of study. In short, our survey of the literature reveals several broad trends and patterns. Initial film network research for example is primarily focused on individuals, with recent papers adopting an expanded scope including production companies and geographic units. Additionally, we observe a disproportionate emphasis on the US film industry in the foundational literature with this bias only amplified by repeated engagement and cross-citation in subsequent papers. We outline a more detailed account of these observations at the conclusion of this essay.
Our aim in providing these insights is to assist scholars interested in utilizing network methods to explore the film industry, as well as other cultural industries, by demonstrating where existing scholarship has clustered into emergent debates and approaches, and where new research might be directed. In particular, we highlight the research lines that have advanced the least, and identify gaps in the existing literature that future network analyses of the film industry could address.
@@ 3 Survey of the Literature
Survey papers typically adhere to the established structure of their respective fields. However, the literature we review spans multiple disciplines, including film industry studies, business studies, cultural policy studies, sociology, and computational sciences. These disciplines and research fields are structured very differently which complicates the task of conducting a coherent survey. One solution would be to categorize the papers by field or discipline. However, discerning the appropriate disciplinary classification for certain papers is not a straightforward process. Another conventional approach would be to distinguish papers based on different network methods. Our emphasis however was on social network papers evaluating the film industry rather than studies simply using film industry data to explicate or illustrate a network method. Excluding purely method-driven papers made this organisational principle less useful.
To address these challenges, we devised a novel solution involving the classification of papers based on node types. Through systematic examination, we found that scholarship exploring networks of individuals involved in the filmmaking process is often substantively and methodologically different from research focused on connections between films, organizations, events, or geographical units. Although a small number of research papers investigate distinct networks with different node types, to our knowledge there has been no significant study to date that considers the interdependencies among multiple types of nodes within film networks.
@@ 3.1.5 Gatekeeping
Network analysis has proved especially valuable to several authors exploring the importance of gatekeeping as both a commentative and as an employment practice in the film industry. [8] test the hypothesis that industry peers who act as evaluators tend to reinforce dominant social beliefs and norms, and are more likely to acknowledge core members over peripheral network members. Using Hollywood data from 1992 to 2004 they constructed film collaboration networks and analyzed the bipartite affiliation networks by creating an individual-by-movie matrix. The authors also created an “accumulative relational profile” by using moving windows of three-year intervals. Their findings suggest that social relationships play a crucial role in determining how talent is perceived and acknowledged. [9] extend this study by drawing on field theory to hypothesize that peers in the film industry favor creators at the core of the collaboration network, while film critics may prefer those at the periphery. These hypotheses are tested using a dataset of over 12,000 crew and cast members who worked on movies distributed by eight major studios in the United States from 1992 to 2004. The dataset includes information on directors, writers, leading and supporting actors/actresses, editors, cinematographers, and production designers, sourced from IMDb. They use a discrete choice model to analyze award decisions by organizations and estimate the choice set of eligible professionals. Collaboration networks are created based on person-film affiliations and analyzed using a core-periphery algorithm. Node classification as core or periphery is then used as a predictor in the models. Results show that being in the periphery negatively affects peer recognition but has a slight positive association with recognition by critics, controlling for other factors.